At what potential difference must an X-ray tube operate to produce X-rays with a
minimum wavelength of 2A. °
1
Expert's answer
2019-02-17T09:58:08-0500

From the conservation of energy:

eU=hc/λ

U=hc/eλ=(1239.841984 eV∙nm)/(1e∙0.2 nm)=6200 V=6.2 kV